OpenStudy (anonymous):

It takes Pam 45 minutes to drive to work and 60 minutes to drive home from work. Write the ratio of the time Pam spends driving home from work to the time she spends driving to work in three different ways

OpenStudy (lukebluefive):

Three possible ratios would be 45:60, 9:12, and 3:4.

OpenStudy (anonymous):

Three different way? =| Speed 1 = 4d/ 3 Speed 2 = d/ 1 Speeed 1 : Speed 2 = 4d/3 : d Therefore the ratio is 4: 3. Is that the correct answer. By the way, I converted 45 min to 3/4th of an hour and 60 min to 1 hour.

OpenStudy (anonymous):

ok thanks